OpenStudy (timinthor):

Please help! What is the slope of the line 4x – 8y – 16 = 0 in simplest form?

OpenStudy (anonymous):

make it like that y = mx + c. How can you do it? first step is- 8y = 4x - 16. Can you do it further?

OpenStudy (mathmale):

You are to find the slope of the given line. As baljeet16 suggests, you could solve the given equation for y and quickly pick out the slope, m, from your result.

OpenStudy (timinthor):

Im not sure, ill try, one sec

OpenStudy (mathmale):

Please share your work. I could give you much more meaningful feedback if you did that.

OpenStudy (timinthor):

Ohhh so the answer would be 4x then?

OpenStudy (mathmale):

I will gladly comment on your work. I will not comment on "answers."

OpenStudy (anonymous):

Y should not have any coefficient look at this y = mx + c.

OpenStudy (timinthor):

If it is 8y = 4x - 16 then the 4x would be in place of the m in y=mx+c

OpenStudy (mathmale):

Note: Slope m does not contain the variable x. You must find a value for m alone.

OpenStudy (timinthor):

So it would just be 4 then

OpenStudy (mathmale):

Need to see your work, please. 4x – 8y – 16 = 0 Add 8y to both sides of this equation. Share your results.

OpenStudy (anonymous):

y = mx + c 8y = 4x - 16 What is the difference between the two? How can you make the second equation similar to 1.

OpenStudy (timinthor):

4x-8y-16=0 +8y +8y 4x-16=8y But then that moves the 4x out of place so im not sure

OpenStudy (mathmale):

No, you are correct. Now, please divide every term by 8, to isolate y.

OpenStudy (timinthor):

Would you just divide the 8y because the others arent like terms

OpenStudy (timinthor):

4x-y-16+0 would be the equation if so

OpenStudy (mathmale):

Like terms have no bearing here. Start with your 4x-16=8y (which is correct) and divide each of the 3 terms by 8.

OpenStudy (mathmale):

Hint: -16 / 8 = - 2. Do the same thing to the other 2 terms.

OpenStudy (mathmale):

What is 4x divided by 8? 4x / 8 ? simplify. You can do nothing with the x...so leave it alone.

OpenStudy (timinthor):

so then youd have 1/2-2=y ?

OpenStudy (mathmale):

You must keep the x ... I did not mean to imply that you eliminate it. No. Try again.

OpenStudy (timinthor):

so 1/2x-2-y=0

OpenStudy (anonymous):

yes thats right.

OpenStudy (mathmale):

We want to solve for y, remember? 4x = -16=8y divide all terms by 8 to obtain (4/8)x - 2 = y. The coefficient of x is your slope. What is your slope? Call it 'm.' reduce it as much as possible.

OpenStudy (mathmale):

All you need to do is to find m and type out its value. m = ?

OpenStudy (mathmale):

m does not include x; rather, m is the coefficient of the variable x in the most recent equation.

OpenStudy (timinthor):

1/2?

OpenStudy (mathmale):

yes. That's it. m=1/2. the slope is positive one half.

OpenStudy (mathmale):

You must know how to go thru this procedure of solving for y and finding the slope, m.

OpenStudy (timinthor):

Thank you :)

OpenStudy (mathmale):

My pleasure. Bye!

OpenStudy (timinthor):

Bye!